OpenStudy (anonymous):

A rectangular tank with no top made from stainless steel.The volume of the box is 250 cm^3.The base is a square with sides x cm.The material for the bottom will cost RM 2 per cm^2 and the material for the sides will cost RM 4 per cm^2.

OpenStudy (anonymous):

i know that the volume is x^2y = 250. then let y y= 250 / x^2 then how to calculate it's minimum cost??

OpenStudy (anonymous):

hehe...i hope no racist here

OpenStudy (anonymous):

volume = x^2h = 250 let h h = 250 / x^2 area = base + 4 ( lenth x height) area = (x^2) 2 + 4 ( 4x * 250/x) A = 2x^2 + 4000 / x differentiate A' = 0 4x - 4000/ x^2 = 0 x = 10 therefore minimum cost = 2 (10)^2 + 4000/ 10 = 600